OpenStudy (rosales0228):

NEED HELP!! GIVING MEDAL AND FAN!!! You invest $1,600 in an account that pays an interest of rate of 6.5% compound continuously. Calculate the balance of your account after 6 years. Round your answer to the nearest hundredth.

OpenStudy (rosales0228):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

You are now using the formula A = P*e^(r*t)

jimthompson5910 (jim_thompson5910):

P = 1600 (initial balance, ie principal) r = 0.065 (interest rate as a decimal) t = 6 years

jimthompson5910 (jim_thompson5910):

'e' is a constant e = 2.71828182846 it's like pi = 3.14...

jimthompson5910 (jim_thompson5910):

your calculator should have an 'e' button so you don't have to type in 2.71828182846 every time

jimthompson5910 (jim_thompson5910):

also the decimal expansion of 2.71828182846 goes on forever without any known pattern

OpenStudy (rosales0228):

i got 2500.39? but it said it was wrong:(

jimthompson5910 (jim_thompson5910):

A = P*e^(r*t) A = 1600*e^(0.065*6) A = ???

OpenStudy (rosales0228):

i got 2500.39 again

jimthompson5910 (jim_thompson5910):

I'm getting 2,363.16927021256 wolfram alpha gets the same thing http://www.wolframalpha.com/input/?i=1600*e^%280.065*6%29

OpenStudy (rosales0228):

hmmm im wondering what im doing wrong

jimthompson5910 (jim_thompson5910):

how are you typing it in?